OpenStudy (anonymous):

Under a dilation, triangle XYZ where X(1, 3), Y(-2, -4), and Z(5, -2.5) becomes triangle X’Y’Z’ where X’(2.5, 7.5), Y’(-5, -10), and Z’(12.5, -6.25). What is the scale factor for this dilation? A) 2/5 B) -2/5 C) 2.5 D) -2.5

OpenStudy (igreen):

https://www.desmos.com/calculator/3qt1junkeu Here's a graph of it. The red is the original(XYZ) and the blue is the new(X'Y'Z').

OpenStudy (igreen):

Does it get bigger or smaller?

OpenStudy (igreen):

Actually, we can just divide the points. Let's take Point X and point X' and divide. Point X: (1, 3) Point X': (2.5, 7.5) 2.5 / 1 = ? 7.5 / 3 = ?

OpenStudy (igreen):

Can you do that? @Addicto_

OpenStudy (anonymous):

2.5/1 = 2.5 7.5/3 = 2.5

OpenStudy (igreen):

Yes, so 2.5 is the scale factor.

OpenStudy (anonymous):

ok thank you :)